40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Message ID: 2059     Entry time: Fri Jun 8 00:11:27 2018
Author: johannes 
Type: Update 
Category: Lab Work 
Subject: optical setup reloaded 

I have been rebuilding the entire optical setup over the past two weeks to make some fundamental changes and fix some things.

  1. The beams in the AOMs are now much smaller than before, on the order of 250 um waists. This allows for faster fall times for ringdown measurements, and higher bandwidth for the eventual amplitude stabilization.
  2. All mirrors are now Coastline super mirrors, between free-space faradays and cavities there are no more economy lenses, only special order optics. Only exception: one half-waveplate each.
  3. The frontal beat is no longer formed between the rejection ports of the polarizers but rather with a small portion of pickoff light.Visibility of the frontal beat is ~70%
  4. Alignment into the EOMs seems to be better, and the RFAM is not so strongly dependent on temperature anymore. I measured 8e-6 in the west and 1.2e-5 in the east.
  5. Mode-matching has been improved to >92% on both paths, which is the dip in reflected light on resonance.

Still have to modify the transmission path, I found that the beam is much too large on the beat PD, which causes a lot of scattering off the shiny rim.

Attachment 1 is an overview of the new layout and attachment 2 shows the west AOM transfer function from commercial driver modulation input to laser intensity, measured by a PDA10CF. I aligned the AOMs to give maximum deflection while minimizing RFAM at 35 MHz, their mod frequency. The East path ended up with a lot more phase delay this way because i had to move it further from the transducer. Deflection efficiency is in both cases about 80% at AOM saturation.

Attachment 1: CRYO_SETUP.pdf  4.078 MB  Uploaded Fri Jun 8 01:36:27 2018  | Hide | Hide all
CRYO_SETUP.pdf
Attachment 2: aom_response_20180605.pdf  23 kB  Uploaded Fri Jun 8 02:04:10 2018  | Hide | Hide all
aom_response_20180605.pdf
ELOG V3.1.3-